OpenStudy (rainbow_dashie):

1. Ethan needs to order T-shirts for all the members. Each T-shirt costs $15 and the shipping fee for the order is $9.99. (b) Ethan has $500 allocated in the club’s budget for T-shirt orders. How many T-shirts can Ethan order with this amount? Show your work and explain your answer.

OpenStudy (rainbow_dashie):

@AnswerMyQuestions

OpenStudy (anonymous):

So we are supposed to use the equation that we already have to figure this out, correct?

OpenStudy (rainbow_dashie):

Yesh

OpenStudy (anonymous):

Hmm...

OpenStudy (rainbow_dashie):

~taps on chin~

OpenStudy (anonymous):

I don't understand how to get the answer :/

OpenStudy (rainbow_dashie):

@mathmath333

OpenStudy (bookworm14):

So the shipping fee for the whole order is about 10$ now we need to subtract 500- 10 = 490$ if the shirts are 15$ each, we need to see how many we can buy for $490 490 / 15 = 32.6 so he can buy 32 shirts

OpenStudy (rainbow_dashie):

Ok holdon 1. Ethan needs to order T-shirts for all the members. Each T-shirt costs $15 and the shipping fee for the order is $9.99. (a) Write an equation showing the total cost, y, for an order of x T-shirts. (b) Ethan has $500 allocated in the club’s budget for T-shirt orders. How many T-shirts can Ethan order with this amount? Show your work and explain your answer. (c) There are 45 members in the club. How much money would Ethan need in order to purchase each member 1 T-shirt? Show your work. Whole question

OpenStudy (rainbow_dashie):

@Bookworm14 Thx!

OpenStudy (bookworm14):

welcome! do you need help with any of the other ones? :)

OpenStudy (rainbow_dashie):

Yes hold on.

OpenStudy (rainbow_dashie):

Its (C) ^

OpenStudy (bookworm14):

ok one sec...

OpenStudy (bookworm14):

If there are 45 members, and each shirt costs 15$, you simply need to multiply. 45 * 15= 675$ Also, i'm not sure if it wants you to add the shipping fee, but if so you just need to add. $675 + $10 = $685 :)

OpenStudy (rainbow_dashie):

Ok thx hold on i have more

OpenStudy (rainbow_dashie):

http://prntscr.com/61gvgu

OpenStudy (rainbow_dashie):

@Bookworm14

OpenStudy (bookworm14):

sorry i got booted out, dang internet lol

OpenStudy (bookworm14):

okay, i know how to do this, just give me a sec to wrk it out and type it in :)

OpenStudy (bookworm14):

Here is all the work Organic Treats and Baked chips both had profits over $350